Question 100)

Professional basketball has truly become a sport that generates interest among fans around the world. More and more players come from outside the United States to play in the National Basketball Association (NBA). You want to develop a regression model to predict the number of wins achieved by each NBA team, based on field goal (shots made) percentage and three-point field goal percentage.

1) Estimate the following multiple regression model (copy and paste the results from either Minitab

Wini=b0+b1FieldGoali+b2Threepoint+eizUjw0L5FgxnrwLrLpt9RP4DhAbTH72jMsYAAAAAS

2) Interpret the meanings of the coefficients of the regression model.

3) Predict the mean number of wins for a team that has a field goal percentage of 45% and a three-point field goal percentage of 35%.

4) Perform a residual analysis on your results and determine whether the regression assumptions are valid.

5) Using t-test, show if the field goal percent or the three point field goal percent are significant at the 0.05 level of significance using the following hypothesis? (Use the p-values to explain without using t-table)

a. Ho: β1=0, Ha:β1≠0 bIogdSYsDmUeNetjQJalhqRQU5Sb5MvNHwE708Fa

b. Ho: β2=0, Ha:β2≠0 N+0gAAAABJRU5ErkJggg==

7) Interpret the meaning of the coefficient of multiple determination (r2) in this problem.

8) Show how to calculate the adjusted r2. Explain the difference between r2 and adjusted r2.

9) At the 0.05 level of significance, perform a F test to determine whether the regression model has any significant variable. Setup your hypothesis and perform the test using p-value to explain.

10) Show how to compute the 95% confidence interval of the estimated coefficients and explain their meanings.

AS THE QUESTION WAS GETTING TOO LONG TO POST, PLEASE FIND A COPYABLE TABLE POSTED UNDER THE NAME "QUESTION 100"
Answer #1

A required regression model to predict the number of wins achieved by each NBA team, based on field goal (shots made) percentage and three-point field goal percentage is given as below:

Regression Statistics

Multiple R

0.692945243

R Square

0.48017311

Adjusted R Square

0.441667414

Standard Error

9.132542725

Observations

30

ANOVA

df

SS

MS

F

Significance F

Regression

2

2080.109911

1040.054956

12.47018402

0.000145913

Residual

27

2251.890089

83.40333662

Total

29

4332

Coefficients

Standard Error

t Stat

P-value

Lower 95%

Upper 95%

Intercept

-263.4928464

63.09278893

-4.176275147

0.000277207

-392.9485546

-134.0371381

Field Goal %

4.409542804

1.242684914

3.548399723

0.001442028

1.859764005

6.959321604

Three-Point Field Goal %

2.806937603

1.656982407

1.694005676

0.101771186

-0.592909427

6.206784632

5)
p-value of Fied Goal (b1) = 0.00144 < alpha
hence it is significant

whereas p-value of Three-point fireld goal(b2) is 0.1018 > alpha
hence it is not significant

7)

r^2 = 0.4802

this means 48.02 % of variation in the number of wins can be explained by this model

8)
adjusted R^2 = 0.4417
= 1 - (1-R^2)(n-1)/(n-k-1)
= 1 - (1- 0.480173)*29/27
= 0.441667

r^2 always increases when we increase more independent variables
but adjusted r^2 penalizes for adding more independent variables
hence for multiple regression, adjusted r^2 is more useful

9)

From this regression model, the p-value or overall model is given as 0.0001459 which is very less than alpha value 0.05,
so we reject the null hypothesis that the given regression model is not statistically significant.
There is sufficient evidence to conclude that this regression model is statistically significant for prediction of the wins for a team.

Please rate if helpful

By HomeworkLib answering guidelines, we have to answer only first 4 sub-parts in multiple sub-parts

Please post rest questions again if left